OpenStudy (anonymous):

Find the exact values of cos 150º and sin 150º.

OpenStudy (anonymous):

Cos150= .6992508065 Sin150= -.7148764296

OpenStudy (anonymous):

here i the answers Cos = –1/2, sin = (3/2 Cos = 1/2, sin = (3/2 Cos = –(3/2, sin = 1/2 Cos =(3/2 , sin = –1/2

OpenStudy (shamim):

\[\cos 150= \cos (180-30)=-\cos30=-\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

i think u can try for sin 150 in this way

OpenStudy (anonymous):

\[would \it be 2. \cos =1/2, \sin= \sqrt{3}/2\]